OpenStudy (anonymous):

Simplify... 2 to the power of 5/2 to the power of 2 the"/" is a fraction..

OpenStudy (anonymous):

2^5=32

jhonyy9 (jhonyy9):

(2^(5/2) )^2 =2^(2*5/2) =2^5 =32

OpenStudy (anonymous):

\[(2^(5/2))^ 2 = 2 ^ ((5/2)*2) = 2^5= 32\]

OpenStudy (campbell_st):

Use the chain rule for powers.... (2^5/2)^2 multiply the powers... 5/2 x 2 = 5 hence 2^5

OpenStudy (anonymous):

Thxs guys.!(: <3